In Exercises 13-20, an equation and two ordered pairs are given. Show that each pair is a solution of the equation. Then graph the two pairs to determine another solution. Answers may vary.
6x − 3y = 3; (1,1), (−1,-3)

Answer by mananth(16946)

6x-3y=3
(1,1)
6*1-3*=3
3=3
.......
& (-1,-3)
6*-1-3*-3=3
-6+9 =3
3=3
Both are solutions to the equation.
6x-3y=3
x=3 y=5 (3,5) is another solution.